Hi. I'm having some difficulty with this problem, Can you help?
A function is given below. Determine the average rate of change of the function between x = 0 and x = 0 + h.
f(t) = √-3t
thanks!

Answer by stanbon(75887)

Determine the average rate of change of the function between x = 0 and x = 0 + h.
f(t) = √-3t
------------------
f(h) = sqrt(-3*h) = sqrt(-3h)
f(0) = sqrt(-3*0) = 0
------------------------------
average = [f(h)-f(0)]/[h-0] = [sqrt(-3h) - 0]/[h-0]
[sqrt(-3h)]/h
======================
Cheers,
Stan H.